Circular No. 5708

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

α ORIONIS

E. Guinan, B. Deeney, and K. Miller report: “Photoelectric photometry being carried out at Villanova University Observatory shows that α Ori is undergoing a rapid decrease in brightness. Beginning in 1992 October the star’s brightness has decreased ∼ 0.45 mag, to V ≃ +0.87 in early 1993 February. The last time α Ori was this faint occurred ∼ 4-4.5 yr ago. We encourage complementary observations over the next few months.”

GRB 930131 W. Wenzel and J. Greiner, Sonneberg Observatory, report that they

have made blink comparison of pairs of plates taken Feb. 3.2 and 4.2 UT of a 2^{o} radius surrounding the γ-ray burst reported on IAUC 5702. No suspicious object was found down to threshold mag ∼ 15.

2CG 135+01 The Compton Observatory EGRET Team (C. von Montigny, K. Bra-

zier, G. Kanbach, and H. A. Mayer-Hasselwander, Max-Planck-Institut für Extraterrestrische Physik; D. L. Bertsch, B. L. Dingus, C. E. Fichtel, R. C. Hartman, S. D. Hunter, J. R. Mattox, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; J. Chiang, J. Fierro, Y. C. Lin, P. F. Michelson, and P. L. Nolan, Stanford University; and E. Schneid, Grumman Aerospace Corporation) reports the detection of the COS-B source 2CG 135+01. The EGRET position excludes the quasar 0241+622 from being the counterpart, but is consistent with the variable radio source GT 0236+610 (Gregory et al. 1989, Ap.J. 339, 1054). No flux variation is evident, but the observation (1991 Nov. 28-Dec. 12) does not cover the phase of the radio outburst. Additional EGRET exposure is planned for the period 1993 Feb. 25-Mar. 9, which corresponds to the phase interval 0.26-0.71 of the 26.5-day period. The radio outburst normally occurs at phase 0.5-0.7. Simultaneous observations at other wavelengths would be of interest. Those planning observations should contact Carl Fichtel at fichtel@lheavx.gsfc.nasa.gov.

PERIODIC COMET SCHWASSMANN-WACHMANN 1 Total V CCD magnitude estimates by H. Mikuz, Ljubljana, Slovenia,

using 0.2-m cameras: Jan. 25.94 UT, [17.7; Feb. 8.76, 12.8.

1993 February 13 Daniel W. E. Green